### v2

# LPR: Large Language Models-Aided Program Reduction

Abstract.
Program reduction is a widely used technique to facilitate debugging compilers by automatically minimizing programs that trigger compiler bugs. Existing program reduction techniques are either generic to a wide range of languages (such as Perses and Vulcan) or specifically optimized for one certain language by exploiting language-specific knowledge (e.g., C-Reduce). However, synergistically combining both generality across languages and optimality to a specific language in program reduction is yet to be explored.
This paper proposes LPR, the first LLMs-aided technique leveraging LLMs to perform language-specific program reduction for multiple languages. The key insight is to utilize both the language generality of program reducers such as Perses and the language-specific semantics learned by LLMs. Concretely, language-generic program reducers can efficiently reduce programs into a small size that is suitable for LLMs to process; LLMs can effectively transform programs via the learned semantics to create new reduction opportunities for the language-generic program reducers to further reduce the programs.
Our thorough evaluation on 50 benchmarks across three programming languages (i.e., C, Rust and JavaScript) has demonstrated LPR’s practicality and superiority over Vulcan, the state-of-the-art language-generic program reducer. For effectiveness, LPR surpasses Vulcan by producing 24.93%, 4.47%, and 11.71% smaller programs on benchmarks in C, Rust and JavaScript, separately. Moreover, LPR and Vulcan have the potential to complement each other. For the C language for which C-Reduce is optimized, by applying Vulcan to the output produced by LPR, we can attain program sizes that are on par with those achieved by C-Reduce. For efficiency, LPR is more efficient when reducing large and complex programs, taking %, %, % less time than Vulcan to finish all the benchmarks in C, Rust and JavaScript, separately.
1. Introduction
Program reduction techniques (Sun et al., 2018; Xu et al., 2023; Regehr et al., 2012; Kalhauge and Palsberg, 2019, 2021; Niemetz and Biere, 2013; Tian et al., 2023b; Wang et al., 2021; Misherghi and Su, 2006; Donaldson et al., 2021) aim to facilitate compiler debugging by minimizing the bug-triggering programs with efficacy and efficiency. Given a program and a property that preserves, program reduction techniques (a.k.a., program reducers) produce a minimal program that still preserves . Program reduction has been widely used in various software engineering tasks (Donaldson and MacIver, 2021), especially in compiler testing and debugging (Livinskii et al., 2020; Lecoeur et al., 2023).
However, a critical challenge in program reduction has not been properly addressed, i.e., the trade-off between generality across languages and specificity to a certain language. Currently, there are two categories of program reduction techniques: language-specific (Regehr et al., 2012; Kalhauge and Palsberg, 2019, 2021) and language-generic (Zeller and Hildebrandt, 2002; Misherghi and Su, 2006; Sun et al., 2018; Xu et al., 2023; Zhang et al., 2023). The former category leverages language-specific semantics to transform and shrink programs in certain languages, while the latter only uses transformations applicable to any programming language. Although language-specific reducers are usually more effective in reduction, designing an effective reducer for a specific language, especially designing language-specific transformations, requires a deep understanding of language features and a significant amount of time and engineering effort. Therefore, only a limited set of languages have custom-designed reducers, such as C (Regehr et al., 2012), Java (Kalhauge and Palsberg, 2019, 2021), and SMT-LIBv2 (Niemetz and Biere, 2013). Meanwhile, language-generic reducers can be applied to diverse languages, but lack the knowledge of language features and semantics and thus are incapable of performing language-specific transformations (e.g., function inlining) that can enable further reduction. As a result, they cannot utilize peculiar features of each language to achieve optimal reduced programs.
This study strives to find a sweetspot between generality across languages and specificity to a certain language, by synergistically combining the strengths of both categories of program reduction techniques. Specifically, we notice that the major limitation of language-generic reducers lies in their incapability to perform language-specific transformations. Language-generic reducers such as Perses stand out as high generality when reducing programs across various programming languages, while they lack awareness of semantic information to achieve further progress. If we could help language-generic reducers conquer this limitation, they are likely to produce smaller results.
Meanwhile, we also notice that recent progress in Large Language Models (LLMs) could be a powerful assistant in performing language-specific transformations, like its performance in other scenarios of software engineering tasks such as code generation and test generation (Sun et al., 2023; Jain et al., 2022; Wei et al., 2023; Deng et al., 2023a; Fan et al., 2023; Xia et al., 2023b; Gu, 2023). Specifically, LLMs are trained with massive programs, and they have started to exhibit the ability to analyze and transform programs in prevalent languages. If we can properly leverage this ability for program reduction, we may have a language-generic reducer being aware of the semantics of various prevalent languages. Besides, the utilization of LLMs can simplify the customization and extension of reducers, as it would be time-consuming and labor-intensive to manually implement a language-specific reducer or add functionality to an existing one (such as C-Reduce which uses the Clang frontend to implement C-specific program transformations).
Challenges of Using LLMs. LLMs are not the silver bullet to program reduction. They face challenges like understanding subtle differences (Li et al., 2023) in code, getting easily distracted by irrelevant context (Shi et al., 2023) and performing poorly with large inputs due to catastrophic forgetting (Kemker et al., 2018; Arora et al., 2019; Goodfellow et al., 2013). Specifically, in program reduction, LLMs may not be capable of reducing large programs directly because of distraction and catastrophic forgetting. Without effective guidance, LLMs are unclear about what transformations to perform.
LLMs-Aided Program Reduction (LPR). We propose LPR (Large language model aided program reduction) in this paper, which is, to the best of our knowledge, the first approach that integrates LLMs for program reduction task. LPR synergistically leverages the strengths of both language-generic program reducers and LLMs. Specifically, LPR alternates between invoking a language-generic reducer (we use Perses in experiments) and the LLM. Initially, Perses efficiently reduces large programs to a size manageable for the LLM. Subsequently, the LLM further transforms Perses’s output based on specific user-defined prompts that dictate the required transformations. Following this, Perses is re-invoked, as transformations made by the LLM often create additional opportunities for reduction. This process iterates until the program cannot be further minimized. For transformations, we have identified five language-generic transformations to enable further reduction: Function Inlining, Loop Unrolling, Data Type Elimination, Data Type Simplification, and Variable Elimination.
To address the aforementioned challenge of using LLMs, LPR is designed with a multi-level prompting approach. In detail, LPR initially requests the LLM to identify a list of potential targets for a given transformation, and then sequentially instructs the LLM to apply the transformation on each target. The multi-level prompt guides the LLM in a more concentrated way, by excluding irrelevant context and other targets that may distract the LLM.
We have conducted extensive evaluations on LPR, illustrating its superiority over Vulcan, the state-of-the-art language-generic algorithm. On three benchmark suites, i.e., Benchmark-C, Benchmark-Rust and Benchmark-JS, LPR produces significantly smaller programs than Vulcan by 24.93%, 4.47% and 11.71%, separately. Moreover, LPR and Vulcan complement each other to some extent. For the C language which C-Reduce is optimized for, by applying Vulcan to the output produced by LPR, we can attain program sizes that are on par with those achieved by C-Reduce. For efficiency, LPR performs comparably to Vulcan. In terms of execution time, LPR is more efficient than Vulcan on reducing complex programs. Furthermore, our detailed analysis indicates that each of the proposed transformations plays a crucial role in the reduction process. We also compare the performance with the multi-level prompt against that without it, illustrating the efficacy of our proposed multi-level prompting approach.
Contribution. This study makes the following contributions.
We introduce LLMs-Aided Program Reduction (LPR), marking the first algorithm that integrates LLMs into the program reduction process. By synergizing the capabilities of both language-generic tools and LLMs, LPR achieves a balance between generality across various languages and awareness of semantics in specific languages. Simultaneously, it maintains flexibility in the design and extension of new transformations.
We propose a multi-level prompting approach to guide LLMs to execute program transformations, and demonstrate its effectiveness in practice. We propose five general-purposed transformations for LLMs to reduce programs or expose more reduction opportunities.
We comprehensively evaluated LPR on 50 benchmarks across three commonly used languages: C, Rust and JavaScript. Results demonstrate LPR’s strong effectiveness and generality.
2. Background
2.1. Program Reduction
Given a program with a certain property, e.g., triggering a compiler bug, the goal of program reduction is to search for a minimal program that still triggers the bug. Program reduction has demonstrated its significant usefulness in removing bug-irrelevant code snippets. The original bug-triggering code (Yang et al., 2011; Le et al., 2014; Sun et al., 2016a; Livinskii et al., 2020) may have thousands of lines, whereas the distilled version from program reduction tools only contains dozens of lines of code (Sun et al., 2016b). Some algorithms are designed to generalize across multiple programming languages, while others are customized for certain languages.
2.1.1. Language-Generic Reducers
Some reducers can be generalized to multiple languages. For instance, given the formal syntax of a programming language, algorithms like HDD and Perses can be used to reduce programs corresponding to that language. HDD parses the language into a parse tree and then applies the DDMin (Zeller and Hildebrandt, 2002) at each level of the tree to remove unnecessary tree nodes as much as possible. Perses goes further than HDD by performing certain transformations on the formal syntax to avoid generating syntactically incorrect program variants. Vulcan extends Perses, by introducing novel auxiliary reducers to exhaustively search for smaller variants by replacing identifiers/sub-trees and deleting local combinations of tree nodes on the parse tree.
However, different languages possess unique semantic features. Although the aforementioned algorithms are relatively efficient (Sun et al., 2018), they are incapable of utilizing unique semantics of a particular language to further reduce a program. For example, these algorithms lacks the ability to perform transformations like function inlining. Although Vulcan can identify more reduction opportunities through transformations such as identifier replacement and local exhaustive search, its approach is akin to ”brute force” enumeration. This method lacks awareness of the given program’s semantics, making it less effective and efficient overall.
2.1.2. Language-Specific Reducers
Previous work has introduced reducers customized for some specific languages. For example, C-Reduce (Regehr et al., 2012) is the most effective reduction tool for C code. It comprises multiple passes that transform the program based on features of the language, thereby making it smaller. Language-specific reducers often rely on static program analysis tools for analysis and modification, e.g., LibTooling (LLVM, 2000) is employed in C-Reduce.
However, developing language-specific reducer is nontrivial. To the best of our knowledge, only a few languages have specific reducers, such as C (Regehr et al., 2012), Java (Kalhauge and Palsberg, 2019, 2021), and SMT-LIBv2 (Niemetz and Biere, 2013). The reason is that the process of designing a reducer for a language, or adding new transformations to an existing reducer, is a time-consuming and labor-intensive task. For instance, in version 2.10.0 of C-Reduce (Regehr, 2012), function inlining was implemented with 604 lines of C++ code. Such challenges impede the development and maintenance of language-specific reducers.
2.2. Large Language Models
Large Language Models refer to a type of deep learning technique that is trained on massively huge data sets for diverse tasks. The advent of LLMs has opened up numerous potential opportunities across diverse research fields. LLMs are not only proficient in processing natural languages but also exhibit substantial capabilities in understanding and processing programming languages. This highlights the promising future and evolutionary prospects in the realm of software engineering. Recently, LLMs have been applied and assessed on various software engineering tasks, such as automatic program repair (Huang et al., 2023; Xia and Zhang, 2023; Xia et al., 2023b, a; Fan et al., 2023) and program generation (Liu et al., 2023; Zhong and Wang, 2023; Tian et al., 2023a).
However, despite the usefulness of LLMs, some researchers (Li et al., 2023) illustrate that current LLMs are weak in distinguishing nuances between programs. Moreover, the memorizing and processing capacity of LLMs deteriorates as the input size grows, a.k.a., catastrophic forgetting (Kemker et al., 2018; Arora et al., 2019; Goodfellow et al., 2013). Moreover, one cannot expect LLMs to automatically complete complex tasks; they must be guided accordingly. Therefore, for program reduction, directly asking LLMs to reduce programs with tens of thousands of lines is impractical.
3. Approach
In this section, we first introduce a motivating example, then we provide an overview of the LPR workflow. We also outline the details of prompts and proposed transformations in the workflow that enable the LLM to function effectively with given programs.
3.1. Motivation
A motivating example is displayed in Figure 1, the original code contains highly nested loops, shown in Figure 0(a). From Figure 0(b) to Figure 0(c), the nested loops are fully unrolled into hundreds of lines via Loop Unrolling, based on the semantic transformations from the LLM. Despite the temporary size increase, the following Perses effectively eliminates all lines except for the bug-relevant one. This is also the final result of LPR, presented in Figure 0(d). By contrast, as shown in Figure 0(e), Vulcan is incapable of escaping the local minima by exhaustively replacing identifiers and tree nodes. Moreover, C-Reduce is not integrated with transformations to unroll loops, and thus cannot fully break down the loop structures. Even though loop unrolling techniques can be added into C-Reduce in future versions, it will be labor-intensive to implement a specific transformation compared to user-defined prompts in natural language.
3.2. Workflow
The overview of the workflow is outlined in Figure 2. Given a bug-triggering program as input, LPR invokes a language-generic reducer and the LLM alternately, until the target program cannot be further reduced. In each iteration, the language-generic reducer efficiently reduces the given program to 1-tree-minimality in syntax level. By contrast, the LLM leverages the semantic knowledge of the language and transforms the program given by the language-generic reducer. This process is guided by user-defined prompts, aiming to expose more reduction potentials to the language-generic reducer.
Algorithm 1 shows LPR’s reduction algorithm. Given as inputs (1) a program targeted for reduction, (2) a property that must be preserved, and (3) the pre-defined prompts, LPR generates a reduced program . prompts contains primaryQuestion and followupQuestion. primaryQuestion instructs the LLM to identify a list of targets to be transformed, and followupQuestion guides the LLM to apply transformation on each target individually. They will be further introduced in section 3.4. Initially, LPR loads a sequence of transformations as delineated in Algorithm 1. It then iterates through each transformation, as detailed from Algorithm 1 to Algorithm 1. section 3.3 displays the details of each transformation.
During this process, for each transformation, the algorithm retrieves a predefined primary question along with a follow-up question on Algorithm 1 – Algorithm 1. LPR first asks the LLM the primary question under the current program. This query aims to guide the LLM to generate a list of specific targets upon which the transformation will be executed. For instance, for Loop Unrolling in the motivating example Figure 0(b), the LLM is asked to identify a list of loops in the given program to be unrolled in the primaryQuestion, and returns a target list .
On Algorithm 1 to 1, LPR uses followupQuestion to guide LLM to apply the transformation on each identified target within the program. In the motivating example, the followupQuestion can be framed as “Given the program { PROGRAM } and the loop for(i=0;...), optimize it via loop unrolling”. The modified program is then extracted from the LLM’s response text on Algorithm 1. In the example, all loops are unrolled into repeated lines of code in Figure 0(c).
In experimental scenarios, given an input program and the prompt, the LLM may generate multiple transformed programs, as the number of responses can be customized. Among all transformed programs, LPR keeps the smallest one that still passes the property test, and discards others. If no transformed program returned from this query satisfies the property, LPR keeps the original one before this query. After each transformation is completed, language-agnostic reducers such as Perses (Sun et al., 2018) are employed to seek additional reduction opportunities, considering that the transformation might have introduced new potentials for further simplification. The algorithm persists in the outermost loop until it reaches a fixpoint, signifying that the program size can no longer be reduced.
3.3. Transformations
To further search for more reduction opportunities on a bug-triggering program via the LLM, we propose five general transformations to guide the LLM, i.e., Function Inlining, Loop Unrolling, Data Type Elimination, Data Type Simplification and Variable Elimination.
Function Inlining. This transformation identifies a function and performs function inlining to eliminate all call sites of this function, and instead substitutes them with corresponding the function body. As functions are prevalent presented in bug-triggering programs, there is significant room for function inlining to reduce tokens or provide further reduction opportunities.
Loop Unrolling. Loop unrolling, also known as loop unwinding, is a widely used loop transformation approach to optimize the execution. In this task, it can also be employed to find more reduction opportunities. Loop Unrolling identifies a loop structure and attempts to unroll the loop into a code snippet repeating a single iteration. Motivation lies in that language-generic reducers may be incapable of dismantling or directly removing the loop structure, while they may be able to reduce the repeated code after loop unrolling, as shown in Figure 1.
Data Type Elimination. Some data types in bug-triggering programs may be irrelevant to the bug, such as identifiers defined by typedef in C, and type alias created by type keyword in Rust. We propose Data Type Elimination to eliminate the alias and replace the occurrence of each alias with its associated original data type.
Data Type Simplification. In programs with complex data types, such as structures, arrays, and pointers, not all components are essential for maintaining bug-triggering properties. For example, a bug-triggering program containing a struct with three integer members can be simplified into three distinct integer variables, and possibly only one variable is essential. To facilitate this simplification, we introduce Data Type Simplification, a strategy designed to transform variables of complex data types into variables of primitive data types, like integers or floats.
Variable Elimination. Intermediate variables are pervasive in programs, and reducing them is desirable in program reduction tasks. Besides, some variables, although not being used, are hard to eliminate. For instance, to remove an unused parameter, both the parameter defined in the function and its corresponding argument passed to the call site of this function should be removed simultaneously. This is hard or even impossible for language-generic reduction tools. Therefore, we propose Variable Elimination to optimize out both intermediate and unused variables.
The proposed transformations are universally applicable across various programming languages, offering a broad utility. By performing these transformations on programs via the LLM, substantial human effort is saved from designing and implementing reducers that target these transformations. While certain existing language-specific reducers like C-Reduce, may already incorporate some of these transformations, e.g., Function Inlining and Variable Elimination, creating new transformation passes remains a non-trivial task for users. Our approach not only simplifies this process but also extends its reach across multiple programming languages.
3.4. Multi-level Prompts
Prompts enable LLMs to apply the transformations mentioned above. We take Function Inlining as an example. We avoid directly instructing the LLM to perform transformations exhaustively, such as inlining all functions in a program in a single query, which might overwhelm its processing capabilities, especially for programs with multiple functions. Instead, we employ a multi-level prompting approach. Figure 3 presents an example of Function Inlining. First, we pose a primary question to the LLM (step ): “Given the following program { PROGRAM }, identify all functions that can be inlined.” Based on the list provided by the LLM (step ), we then ask a series of follow-up questions (step and step ) like “Given the following program { PROGRAM } and the specified function { fn1 }, optimize { fn1 } out via function inlining.”, and the LLM do the transformations accordingly (step and step ). This strategy excludes irrelevant context and ensures that the queries are more targeted, thereby increasing the likelihood of the LLM generating high-quality results. For other transformations, the prompts follow a similar template — first prompting the LLM to identify a target list, and then instructing it to attempt optimization of each target.
4. Evaluation
In this section, we evaluate the effectiveness and efficiency of LPR. Specifically, we conducted the following research questions.
RQ1. What is the effectiveness of LPR in program reduction?
RQ2. What is the efficiency of LPR in program reduction?
RQ3. What is the effectiveness of each transformation in LPR?
4.1. Experimental Setup
Within the workflow of LPR, we employ Perses (Sun et al., 2018) as the language-agnostic reducer due to its superior efficiency compared to Vulcan. Additionally, we utilize OpenAI API (ope, 2023a), specifically the gpt-3.5-turbo-0613 version, to serve as the LLM. We also develop a variant named LPR+Vulcan, which invokes Vulcan to further reduce the program after LPR finishes. For a fair comparison, all algorithms were executed in a single-process, single-thread environment.
Benchmarks. To measure the effectiveness and efficiency of LPR across various languages, we employ three benchmark suites: Benchmark-C, Benchmark-Rust and Benchmark-JS. The Benchmark-C, previously collected and utilized by previous studies (Xu et al., 2023; Tian et al., 2023c; Zhang et al., 2023), comprises 20 large complex programs triggering real-world bugs in LLVM or GCC. Benchmark-Rust, incorporating 20 bug-triggering Rust programs, has also been used in prior research (Xu et al., 2023). We further craft Benchmark-JS, a non-public benchmark suite, for this study. Specifically, we use FuzzJIT (Wang et al., 2023) to fuzz a prevalent JavaScript engine, i.e., JavaScriptCore (version c6a5bcc), and then randomly collect 10 programs that cause miscompilations in JIT compiler. Since the programs and reduced programs in Benchmark-JS are not publicly available and thus not in the training sets of LLMs. The evaluation on Benchmark-JS helps us investigate whether LPR suffers from the data leakage problem (Wu et al., 2023). In total, the evaluation benchmarks encompass 50 programs triggering real-world bugs in compilers, spanning across three popular programming languages.
Baselines. In all three benchmark suites, we use Perses and Vulcan as baselines. Perses stands out as a highly effective and efficient program reduction tool. To avoid the occurrence of syntactical invalid variants during the reduction process, it transforms and normalizes the formal syntax of a programming language. Vulcan (Xu et al., 2023), building upon Perses, provides three manually designed auxiliary reducers to further search for reduction opportunities on results from Perses. Compared to Perses, Vulcan achieves a reduction in the number of tokens, albeit at the expense of increased running time. They are both language-generic and are applicable across a broad spectrum of programming languages. We also include C-Reduce (v2.9.0) as an additional baseline. C-Reduce not only stands as the most effective algorithm for C, it can also be applied to other languages, though not customized for them.
Configuration. If not otherwise specified, our experiments are conducted by invoking OpenAI API (version gpt-3.5-turbo-0613), with the proposed multi-level prompt and transformations in section 3.3 and section 3.4. To effectively harness the inherent randomness of LLMs, we set temperature=1.0. This high value encourages the LLM to generate more diverse outcomes (ope, 2023c). Additionally, we employ n=5 to generate five distinct results for every query (ope, 2023b), enabling us to choose the smallest passing program as the optimal result. All the rest configurations are set to their default values.
4.2. RQ1: Effectiveness
We measure the effectiveness of LPR, LPR+Vulcan and baseline algorithms via the final program size in token. A smaller size is favored, as it signifies the removal of more bug-irrelevant code, thereby saving developers more manual effort. The effectiveness of each algorithm on all three benchmark suites is presented in Table 1. Due to the randomness of LPR, we repeat five times for LPR and LPR+Vulcan on every benchmark, and display the mean and standard deviation in the table. On each benchmark, the minimal results are highlighted in bold.
Benchmark | Original | Perses | Vulcan | C-Reduce | LPR | LPR+Vulcan | |||
LLVM-22382 | 21,068 | 144 | 108 | 70 | 73.2 | 1.6 | 69.8 | 1.8 | |
LLVM-22704 | 184,444 | 78 | 62 | 42 | 43.6 | 3.1 | 41.8 | 3.3 | |
LLVM-23309 | 38,647 | 464 | 303 | 118 | 105.8 | 9.3 | 91.2 | 8.2 | |
LLVM-23353 | 30,196 | 98 | 91 | 74 | 68.8 | 8.0 | 66.6 | 6.5 | |
LLVM-25900 | 78,960 | 239 | 104 | 90 | 93.4 | 11.1 | 84 | 6.5 | |
LLVM-26760 | 209,577 | 120 | 56 | 43 | 62.8 | 18.6 | 52.6 | 5.4 | |
LLVM-27137 | 174,538 | 180 | 88 | 50 | 69.2 | 19.2 | 65 | 20.2 | |
LLVM-27747 | 173,840 | 117 | 79 | 68 | 87.8 | 2.5 | 63.2 | 2.2 | |
LLVM-31259 | 48,799 | 406 | 282 | 168 | 184.0 | 51.2 | 114.4 | 10.9 | |
GCC-59903 | 57,581 | 308 | 198 | 105 | 209.8 | 72.1 | 166.4 | 64.0 | |
GCC-60116 | 75,224 | 443 | 247 | 168 | 188.8 | 52.4 | 127.6 | 34.8 | |
GCC-61383 | 32,449 | 272 | 195 | 84 | 113.2 | 13.6 | 105.2 | 5.4 | |
GCC-61917 | 85,359 | 150 | 103 | 65 | 78.4 | 10.6 | 73.4 | 5.5 | |
GCC-64990 | 148,931 | 239 | 203 | 65 | 143.4 | 58.0 | 119 | 50.1 | |
GCC-65383 | 43,942 | 153 | 84 | 72 | 64.2 | 2.7 | 64.2 | 2.7 | |
GCC-66186 | 47,481 | 327 | 226 | 115 | 97.8 | 17.9 | 94.2 | 12.0 | |
GCC-66375 | 65,488 | 440 | 227 | 56 | 56.0 | 5.1 | 56.0 | 5.1 | |
GCC-70127 | 154,816 | 301 | 230 | 84 | 95.0 | 3.8 | 73.6 | 3.3 | |
GCC-70586 | 212,259 | 426 | 223 | 130 | 235.4 | 30.2 | 156.8 | 12.6 | |
GCC-71626 | 6,133 | 51 | 38 | 46 | 44.6 | 3.4 | 36.6 | 0.9 | |
Benchmark-C | Mean | 94,487 | 247.8 | 157.4 | 85.7 | 105.8 | 4.4 | 86.1 | 2.9 |
Rust-44800 | 801 | 467 | 284 | 473 | 124.6 | 32.4 | 118.6 | 34.4 | |
Rust-66851 | 936 | 728 | 713 | 654 | 414.2 | 273.9 | 331.2 | 257.8 | |
Rust-69039 | 190 | 114 | 101 | 110 | 97.2 | 8.0 | 90.8 | 10.9 | |
Rust-77002 | 347 | 263 | 247 | 264 | 96.0 | 27.9 | 96.0 | 27.9 | |
Rust-77320 | 173 | 40 | 40 | 40 | 40.0 | 0.0 | 39.0 | 0.0 | |
Rust-77323 | 81 | 13 | 13 | 13 | 13.0 | 0.0 | 13.0 | 0.0 | |
Rust-77910 | 63 | 34 | 21 | 23 | 29.2 | 2.7 | 21.0 | 0.0 | |
Rust-77919 | 132 | 74 | 62 | 70 | 62.2 | 14.3 | 58.2 | 7.7 | |
Rust-78005 | 182 | 102 | 102 | 75 | 102.0 | 0.0 | 102.0 | 0.0 | |
Rust-78325 | 65 | 29 | 26 | 34 | 29.0 | 0.0 | 26.0 | 0.0 | |
Rust-78651 | 957 | 17 | 9 | 12 | 16.6 | 0.5 | 11.0 | 2.7 | |
Rust-78652 | 263 | 56 | 49 | 49 | 53.6 | 2.2 | 49.0 | 0.0 | |
Rust-78655 | 28 | 26 | 26 | 26 | 26.0 | 0.0 | 26.0 | 0.0 | |
Rust-78720 | 121 | 72 | 56 | 51 | 58.4 | 2.1 | 56.6 | 0.5 | |
Rust-91725 | 513 | 174 | 86 | 101 | 68.6 | 57.6 | 55.2 | 18.4 | |
Rust-99830 | 448 | 299 | 277 | 160 | 271.6 | 12.5 | 230.2 | 62.6 | |
Rust-111502 | 192 | 166 | 157 | 161 | 104.8 | 7.2 | 103.6 | 8.2 | |
Rust-112061 | 556 | 458 | 442 | 450 | 385.0 | 32.6 | 380.4 | 31.8 | |
Rust-112213 | 866 | 736 | 635 | 732 | 653.0 | 34.0 | 618.8 | 22.9 | |
Rust-112526 | 644 | 382 | 338 | 545 | 304.0 | 30.3 | 283.2 | 23.3 | |
Benchmark-Rust | Mean | 378 | 212.5 | 184.2 | 202.2 | 147.5 | 11.1 | 135.5 | 10.1 |
JS-1 | 244 | 52 | 41 | 41 | 25.6 | 0.5 | 24.8 | 1.6 | |
JS-2 | 112 | 51 | 41 | 40 | 34.0 | 6.5 | 27.8 | 3.5 | |
JS-3 | 125 | 57 | 41 | 47 | 51.6 | 12.1 | 35.4 | 7.7 | |
JS-4 | 185 | 65 | 35 | 47 | 33.4 | 2.2 | 28.2 | 6.9 | |
JS-5 | 178 | 66 | 38 | 52 | 41.6 | 2.2 | 33.8 | 5.5 | |
JS-6 | 152 | 57 | 30 | 45 | 20.0 | 2.8 | 19.2 | 2.7 | |
JS-7 | 144 | 46 | 38 | 38 | 34.0 | 0.0 | 27.2 | 6.3 | |
JS-8 | 121 | 55 | 47 | 45 | 40.6 | 6.3 | 33.0 | 17.4 | |
JS-9 | 87 | 50 | 30 | 32 | 23.8 | 2.5 | 18.8 | 5.2 | |
JS-10 | 63 | 56 | 41 | 43 | 34.8 | 5.1 | 27.0 | 6.3 | |
Benchmark-JS | Mean | 141 | 55.5 | 38.2 | 43.0 | 33.9 | 1.6 | 27.5 | 5.9 |
Benchmark-C. On this benchmark suite, Perses reduces the programs to an average of 247.8 tokens. Building upon this, Vulcan further compresses the average program size into 157.4, i.e., thereby continues to decrease the program size by 34.35%. Despite Vulcan’s notable reduction progress, LPR is still capable of continuing to push the limit of Perses, and reduces the programs in Benchmark-C into 105.8 tokens on average across all five runs. It cuts down the average program size of Perses by 51.33%, outperforming Vulcan significantly by 24.93% (proved by a p-value of 0.002). C-Reduce stands out by achieving the lowest average program size, i.e., 85.7 tokens. This performance is anticipated as C-Reduce incorporates various transformation passes specifically designed for C. Despite relying on only general transformations, LPR+Vulcan still achieves performance comparable to C-Reduce, averaging 86.1 tokens. Moreover, it outperforms C-Reduce in 13 out of 20 benchmarks, highlighting its effectiveness with only language-generic transformations.
Benchmark-Rust. On Benchmark-Rust, Perses and Vulcan produce programs with 212.5 and 184.2 tokens on average, separately. LPR and LPR+Vulcan further shrink the average program size into 147.5 and 135.5 tokens. C-Reduce produces the second-largest programs on average, only smaller than Perses. This is anticipated since C-Reduce lacks specialized transformations for Rust.
Further analysis into these benchmarks reveals that Vulcan and transformations in LPR are complementary on Benchmark-Rust. Vulcan shows higher effectiveness on reducing relatively smaller programs, as evidenced by the average original size of 43 tokens in the 9 programs where it is better than LPR. In contrast, the 15 programs where LPR is proved better than Vulcan have an average of 269 tokens, indicating its proficiency on reducing relatively larger programs. Our speculation is that Vulcan and LPR target different reduction opportunities. Vulcan performs identifier/sub-tree replacement and local exhaustive search. Such reducers, while lacking in semantic analysis, find reduction opportunities in a ”brute-force” manner and is particularly effective at uncovering less obvious reduction opportunities. On the other hand, LPR employs more semantic and intelligent transformations, adeptly and systematically analyzing and reducing a complex program step by step. Moreover, results of LPR+Vulcan in the last column prove the complementary characteristic between LPR and Vulcan, which achieve the best in 16 out of the total 20 benchmarks.
Benchmark-JS. Programs in Benchmark-JS are much simpler and smaller than those in the previous two benchmark suites. Therefore, even Perses alone is capable of reducing the programs to only 55.5 tokens. Following this, Vulcan, LPR and LPR+Vulcan achieve 38.2, 33.9 and 27.5 tokens, further reducing the average results by 30.35%, 38.66% and 38.66%, separately. Similar to its performance on Rust, C-Reduce cannot outperform the aforementioned algorithms on JavaScript, due to its lack of employment of JavaScript’s semantics. The evaluation results also serve to demonstrate that the performance exhibited by LPR is not attributable to data leakage. These benchmarks were collected by the authors via fuzzing, and the optimal results remain inaccessible to the public, thereby precluding any possibility of LLMs memorizing them.
4.3. RQ2: Efficiency
In this research question, we measure the time consumed by each technique on three benchmarks. Shorter time indicates higher efficiency. Table 2 shows the results. Since both Vulcan and LPR perform reduction on top of Perses’s results, it is impossible for these two algorithms to take less time than Perses. Besides, as a highly efficient tree-based reduction algorithm, Perses is generally faster than C-Reduce. Therefore, we focus on the comparison among Vulcan, C-Reduce and LPR.
Benchmark | Perses | Vulcan | C-Reduce | LPR | LPR+Vulcan | |||
LLVM-22382 | 0:06:46 | 0:17:03 | 0:14:46 | 0:39:43 | 0:17:13 | 0:44:33 | 0:16:46 | |
LLVM-22704 | 0:33:58 | 0:38:03 | 0:22:38 | 0:48:53 | 0:01:03 | 0:52:17 | 0:01:23 | |
LLVM-23309 | 0:22:34 | 2:02:39 | 0:48:47 | 1:35:13 | 0:13:36 | 2:05:40 | 0:14:16 | |
LLVM-23353 | 0:10:31 | 0:15:05 | 0:13:36 | 0:25:23 | 0:06:05 | 0:30:47 | 0:06:33 | |
LLVM-25900 | 0:09:53 | 0:23:08 | 0:22:04 | 0:44:03 | 0:07:56 | 0:54:33 | 0:07:21 | |
LLVM-26760 | 0:21:54 | 0:34:23 | 0:32:25 | 0:54:40 | 0:18:08 | 1:04:28 | 0:16:51 | |
LLVM-27137 | 1:54:41 | 3:15:20 | 2:21:43 | 2:33:15 | 0:09:08 | 3:13:24 | 0:08:48 | |
LLVM-27747 | 0:13:32 | 0:28:02 | 0:25:30 | 0:32:04 | 0:02:28 | 0:45:53 | 0:03:17 | |
LLVM-31259 | 0:32:30 | 4:03:54 | 1:13:20 | 4:08:08 | 0:34:26 | 5:01:39 | 0:54:49 | |
GCC-59903 | 0:48:47 | 1:21:15 | 1:23:10 | 2:14:57 | 0:45:14 | 2:44:43 | 0:48:09 | |
GCC-60116 | 0:36:18 | 2:02:01 | 1:15:45 | 3:16:38 | 0:31:53 | 4:25:39 | 0:34:36 | |
GCC-61383 | 0:44:59 | 3:50:40 | 1:00:39 | 2:52:39 | 0:30:41 | 4:36:47 | 0:22:35 | |
GCC-61917 | 0:14:57 | 0:24:16 | 0:44:26 | 0:37:28 | 0:07:37 | 0:43:30 | 0:07:31 | |
GCC-64990 | 0:51:05 | 1:27:12 | 1:20:05 | 2:03:05 | 0:31:38 | 2:22:37 | 0:22:33 | |
GCC-65383 | 0:17:05 | 0:37:02 | 0:33:45 | 0:46:53 | 0:05:25 | 0:59:14 | 0:05:08 | |
GCC-66186 | 0:41:19 | 5:35:16 | 1:24:13 | 2:39:23 | 0:19:26 | 4:05:34 | 0:36:49 | |
GCC-66375 | 0:46:28 | 3:59:57 | 2:02:40 | 2:30:08 | 0:10:24 | 3:06:41 | 0:10:44 | |
GCC-70127 | 0:44:47 | 4:45:15 | 1:40:01 | 2:23:13 | 0:20:36 | 3:24:03 | 0:20:20 | |
GCC-70586 | 1:33:35 | 6:53:31 | 1:37:16 | 6:24:35 | 1:26:57 | 10:36:26 | 2:32:53 | |
GCC-71626 | 0:00:40 | 0:01:18 | 0:04:06 | 0:07:38 | 0:01:27 | 0:08:11 | 0:01:32 | |
Benchmark-C | Mean | 0:35:19 | 2:08:46 | 0:59:03 | 1:54:54 | 0:05:17 | 2:37:20 | 0:08:05 |
Rust-44800 | 0:13:32 | 1:58:31 | 1:33:17 | 1:47:29 | 0:31:37 | 2:15:39 | 0:42:02 | |
Rust-66851 | 0:59:47 | 8:49:11 | 1:32:02 | 11:21:39 | 9:56:51 | 16:43:40 | 12:54:53 | |
Rust-69039 | 0:07:54 | 1:25:33 | 0:10:05 | 0:24:13 | 0:05:33 | 0:39:22 | 0:05:51 | |
Rust-77002 | 0:04:12 | 0:20:17 | 0:29:18 | 0:52:27 | 0:15:54 | 0:58:21 | 0:15:18 | |
Rust-77320 | 0:00:06 | 0:01:22 | 0:01:51 | 0:02:36 | 0:00:32 | 0:04:05 | 0:00:32 | |
Rust-77323 | 0:00:01 | 0:00:11 | 0:00:37 | 0:00:16 | 0:00:03 | 0:00:27 | 0:00:04 | |
Rust-77910 | 0:00:08 | 0:00:47 | 0:01:12 | 0:04:58 | 0:01:34 | 0:05:44 | 0:01:35 | |
Rust-77919 | 0:00:17 | 0:02:46 | 0:05:29 | 0:08:45 | 0:05:08 | 0:11:18 | 0:04:52 | |
Rust-78005 | 0:00:10 | 0:01:57 | 0:02:30 | 0:10:44 | 0:01:32 | 0:12:55 | 0:01:31 | |
Rust-78325 | 0:00:02 | 0:00:28 | 0:01:32 | 0:00:46 | 0:00:35 | 0:01:19 | 0:00:34 | |
Rust-78651 | 0:00:04 | 0:00:23 | 0:01:09 | 0:01:33 | 0:00:33 | 0:02:02 | 0:00:36 | |
Rust-78652 | 0:00:08 | 0:01:32 | 0:03:01 | 0:02:53 | 0:02:02 | 0:04:38 | 0:01:59 | |
Rust-78655 | 0:00:01 | 0:00:49 | 0:01:30 | 0:02:20 | 0:00:31 | 0:03:14 | 0:00:32 | |
Rust-78720 | 0:00:16 | 0:03:47 | 0:06:31 | 0:11:44 | 0:04:47 | 0:13:52 | 0:04:55 | |
Rust-91725 | 0:03:36 | 0:17:48 | 0:37:19 | 0:16:13 | 0:03:55 | 0:23:29 | 0:02:11 | |
Rust-99830 | 0:48:23 | 20:08:32 | 11:12:22 | 4:23:44 | 1:02:03 | 24:28:10 | 1:37:03 | |
Rust-111502 | 0:00:55 | 0:10:35 | 0:10:23 | 0:37:29 | 0:11:08 | 0:44:39 | 0:11:13 | |
Rust-112061 | 0:34:50 | 4:48:04 | 1:15:44 | 5:10:02 | 3:12:51 | 8:01:03 | 3:04:33 | |
Rust-112213 | 0:56:40 | 15:21:26 | 1:08:21 | 7:33:21 | 2:44:52 | 17:07:36 | 3:34:22 | |
Rust-112526 | 0:45:36 | 2:55:07 | 1:35:52 | 3:33:13 | 1:21:21 | 4:52:26 | 1:54:13 | |
Benchmark-Rust | Mean | 0:13:50 | 2:49:27 | 1:00:30 | 1:50:19 | 0:40:46 | 3:51:42 | 0:44:16 |
JS-1 | 0:01:29 | 0:29:19 | 0:06:59 | 0:11:47 | 0:05:01 | 0:14:32 | 0:05:00 | |
JS-2 | 0:02:15 | 0:17:10 | 0:11:47 | 0:14:36 | 0:01:12 | 0:29:51 | 0:04:21 | |
JS-3 | 0:01:29 | 0:26:35 | 0:06:02 | 0:09:00 | 0:00:14 | 0:29:30 | 0:07:58 | |
JS-4 | 0:00:19 | 0:15:43 | 0:01:44 | 0:07:22 | 0:02:35 | 0:37:07 | 0:03:57 | |
JS-5 | 0:00:14 | 0:04:25 | 0:01:39 | 0:10:32 | 0:04:27 | 0:22:49 | 0:04:32 | |
JS-6 | 0:02:59 | 0:23:20 | 0:11:33 | 0:13:24 | 0:07:49 | 0:18:43 | 0:08:08 | |
JS-7 | 0:00:15 | 0:03:30 | 0:16:21 | 0:05:28 | 0:01:06 | 0:10:52 | 0:01:09 | |
JS-8 | 0:00:44 | 0:13:48 | 0:02:21 | 0:08:18 | 0:02:20 | 0:13:04 | 0:02:10 | |
JS-9 | 0:01:32 | 0:14:14 | 0:07:50 | 0:09:09 | 0:02:43 | 0:15:14 | 0:03:40 | |
JS-10 | 0:01:31 | 0:17:44 | 0:07:24 | 0:15:04 | 0:04:13 | 0:32:32 | 0:05:37 | |
Benchmark-JS | Mean | 0:01:17 | 0:16:35 | 0:07:22 | 0:10:28 | 0:01:26 | 0:22:25 | 0:02:33 |
In the Benchmark-C, compared to Vulcan and LPR, C-Reduce generally has a shorter reduction time. This is expected, as C-Reduce’s transformations are specifically designed for C languages, whereas Vulcan approaches the problem in a more unguided and brute-force manner, and LPR’s attempts are not specifically designed for C and rely more on the efficiency of LLMs. On average, LPR takes 1:54:54, which is % shorter than 2:08:46 of Vulcan. However, in terms of the average percentage difference in time, LPR requires 45.83% more time compared to Vulcan in Benchmark-C. The main reason of such a result is that LPR is more efficient than Vulcan when the program is large and complex, as the transformations it performs are aware of the semantics and have a higher success rate in reducing the program. However, when the program is small and simple, Vulcan can finish quickly since its search space becomes considerably small, but for LPR, the time consumed by the LLM is not decreased significantly and becomes dominant, which makes LPR less efficient than Vulcan in these benchmarks. On Benchmark-Rust, The results indicate a similar trend, i.e., LPR tends to be more efficient when reducing complex programs while Vulcan is more efficient in small and simple benchmarks. If we only keep benchmarks where both tools take longer than one hour, LPR requires 4.15% and 21.69% less time compared to Vulcan.
In our analysis of Benchmark-Rust, we observed that both LPR and Vulcan can consume an extreme long time on certain benchmarks. For instance, Vulcan requires 20 hours for Rust-99830, while LPR takes a similar duration on Rust-66851 in a specific run. This extensive time consumption is often due to the frequent invocation of Perses, which only achieves marginal progress with each transformation. The prolonged duration of Perses is primarily attributed to the strict syntax of the Rust language. Considering that program reduction is an NP-complete problem, this inefficiency might be optimized in future, whereas it cannot be completely eliminated.
After further in-depth analysis, another interesting fact emerges. On the three benchmark suites, the average time taken by LPR is 1.915, 1.839, and 0.174 hours, respectively. However, within these durations, the time spent waiting for the LLM responses accounts for %, %, and %. This experiment involves invoking the OpenAI API, and might have been limited by high user traffic and few computational resources allocated. Considering the ongoing advancements in LLMs technology, we believe that LPR’s efficiency will be substantially improved in the future.
4.4. RQ3: Effectiveness of Each Transformation
To answer this question, we delve into the impact of each transformation on program size, alongside their potential to escape the local minimal program and unlock new reduction opportunities. Our in-depth analysis focuses on Benchmark-C because of the complexity of bug-triggering programs.
For each transformation, we measure how the program size change in each benchmark in Benchmark-C after a specific transformation is performed, and plot the size changes into a box-plot and red dots, as shown in Figure 4. Additionally, since each transformation is immediately followed by an invocation of Perses, we also monitor the cumulative size changes resulting from both the transformation and its subsequent Perses reduction. These changes are depicted by the blue dots in the second box-plot of each subplot in Figure 4. Given that a transformation is performed in every iteration, we compute the average size change by dividing the total sum of size changes for that transformation by the number of iterations. This approach enables us to thoroughly comprehend how each transformation influences program size and assess its capacity to provide further reduction opportunities to Perses.
According to size changes induced by each transformation alone, i.e., the left box-plot with red dots, we can find two trends. First, Function Inlining, Data Type Elimination and Variable Elimination are more likely to reduce the program size by themselves, with an average of size change -13.2, -4.7 and -4.3, separately. However, for the rest two transformations, i.e., Loop Unrolling and Data Type Simplification, most of the program sizes increase instead, with an average of 14.9 and 1.3 respectively. This is expected, as such transformations will generally transform the program into a larger one. Loop Unrolling, disassembles a loop into repeated lines of code, and leads to size increase temporarily. Data Type Simplification can dismantle a variable in complex data type, e.g., structure, into a list of members in primary data types, which may requires more tokens to declare and initialize each variable.
For size changes induced by both a transformation and the subsequent execution of Perses, i.e., the right box-plot with blue dots, all of the proposed transformation result in size decreases. The fact that the right box-plot is generally lower than the left one indicates that Perses often further removes tokens after the transformation is applied. For Loop Unrolling and Data Type Simplification, even though they usually introduce more tokens to the program, they expose reduction opportunities for the following execution of Perses, and eventually result in a smaller result.
To further understand the impact of each transformation on the entirety of Benchmark-C, we provide a detailed analysis of their contributions. Specifically, on the 5 repeated experiments on 20 benchmarks, we calculate the average size reduction brought about by each transformation across all these 100 runs by summing up the size decreases attributed to the transformation in each benchmark and then computing the mean. This is illustrated in Figure 5. Additionally, we quantify the prevalence of each transformation by counting the number of benchmarks in which it induces a size decrease, as demonstrated in Figure 6. These metrics together offer a comprehensive view of how each transformation influences file sizes and how frequently they take effect within Benchmark-C.
From Figure 5, it is evident that every transformation contributes to further size reduction in Benchmark-C. Specifically, while Function Inlining is responsible for a reduction of 88.2 tokens on average, contributing 62.12% to the overall decrease, Loop Unrolling shows the minimal impact, contributing only 4.8 tokens, 3.35% to the overall decrease. This highlights the varying degrees of influence each transformation has on code size. Further insights from Figure 6 reveal that Data Type Elimination affects all benchmarks, likely due to the ubiquitous presence of typedef across all benchmarks. In contrast, Loop Unrolling is the least prevalent transformation, affecting merely 6 on average out of 20 benchmarks. This outcome is expected, considering that not all programs involve loop structures, and not every loop is irrelevant to the bugs in compilers. Besides, the relatively higher standard deviations observed in the Data Type Simplification and Loop Unrolling suggest that these transformations present greater challenges for the LLM to manage effectively.
5. Discussion
In this section, we discuss the effectiveness of multi-level prompts and, the performance of the LLM under different temperatures.
5.1. The Effectiveness of Multi-level Prompt
To validate the effectiveness of our proposed multi-level prompt, we design the corresponding single-level prompt and compare its effectiveness against the multi-level prompt. In detail, different from multi-level prompt, single-level prompt merges the primaryQuestion and followupQuestion into a single prompt, e.g., “Given the following program { PROGRAM }, identify one function that can be inlined, and inline it.” for Function Inlining.
In 5 repeated experiments on Benchmark-C, the single-level prompting approach results in tokens, which is far less effective than results from the multi-level prompting approach, i.e., tokens. Our explanation is that, even though such a single-level prompt is more compact, it makes LLMs less concentrated on a specific target, and thus LLMs may omit some targets.
5.2. The Impact of Temperature
In our experiments, we consistently set the temperature parameter of LLMs to 1.0. To measure how this parameter affects the performance, we rerun experiments under multiple temperatures, i.e., 0.75, 0.5, 0.25, 0. Note that higher temperature instructs the LLM to generate more creative and diverse results. Due to limited resources and time, we evaluate on 10 benchmarks in Benchmark-C that demonstrated the fastest completion times under the default configuration.
Perses | Vulcan | LPR with gpt-3.5-turbo-0613 | |||||
Mean | 161.4 | 102.8 |
As shown in Table 3, the performances under most of the temperatures are similar, while the exception is t=0, with the average size worse than others. According to the documentation of temperature (ope, 2023c), t=0 will actually use a small threshold above 0. Our speculation is that a low temperature restrains the diversity of outputs, impeding LPR exploring local minima in different runs, which is helpful in program reduction tasks.
Given that program reduction is an NP-complete problem, randomness of LLMs has its advantages and drawbacks. Take the results of RQ1 in Table 1 as an example. On the one hand, randomness allows LPR to explore more distinct local minima, and sometimes generates smaller programs than C-Reduce in five repeated experiments on each benchmark. On the other hand, randomness of LLMs introduces variability. While the standard deviation remains below 10 in most benchmarks, it may significantly increase in certain benchmarks, such as in GCC-59903. This variation can be attributed to the high complexity of the given program. In such scenarios, LLMs might not consistently execute accurate transformations, resulting in a range of local minima and divergent results.
5.3. Threats to Validity
In this section, we discuss potential factors that may undermine the validity of our experimental results.
5.3.1. Threats to Internal Validity
The main internal threat comes from the potential data leakage problem. That is, do LLMs provide reasonable transformation through step-by-step analysis, or just simply memorize the minimal programs for the benchmark suites, which may be publicly available on the internet?
We mitigate this threat from several perspectives. First, program reduction is a task involving programs distinct from those used in program repair or program synthesis tasks. For instance, LLMs can learn from large datasets about code generation. On the contrary, programs in program reduction tasks are generally large, complex, and most importantly, randomly generated to trigger compiler bugs, with no other specific purpose. Their random and chaotic characteristics make it highly unlikely for LLMs to memorize such disorganized content, thus reducing the risk of data leakage.
Second, we asked LLMs to verify whether they remember any bug-triggering programs from public bug-tracking systems. For instance, we ask LLMs “Do you know the minimal bug-triggering program in GCC’s bugzilla, with id=66186?” and the results confirm that the LLMs do not retain the memory of these programs. Moreover, even in scenarios where the LLMs might coincidentally memorize certain minimal programs, it is improbable for it to link a random-looking, featureless code snippet with a specific memorized program, especially when no explicit bug ID is provided.
Furthermore, Benchmark-JS, one of the benchmark suites used, was created using JIT fuzzing tools by the authors and is not publicly accessible. This exclusivity ensures that the LLMs’ performance on these benchmarks reflects their ability to handle unseen and novel programs, thereby showcasing their effectiveness in managing new challenges without relying on memorized data. This approach significantly mitigates the risk of data leakage and demonstrates LLMs’ capacity for genuine problem-solving and analysis.
5.3.2. Threats to External Validity
One threat to external validity is the generality of LPR across languages. We evaluate LPR on three prevalent programming languages, namely, C, Rust and JavaScript. We believe these benchmark suites have demonstrated LPR’s high generality on diverse programs. However, without further experiments on a particular language, we still cannot ensure the effectiveness and efficiency of LPR on that language.
An additional threat is the applicability of our approach across different LLMs. To mitigate this threat, we repeat the experiments with other versions of ChatGPT, i.e., gpt-3.5-turbo-1106 and gpt-3.5-turbo-16k-0613. The experimental results show no significant differences from results on gpt-3.5-turbo-0613. Furthermore, as LLMs continue to evolve, we anticipate improvements in both quality and time of code processing.
6. Related Work
We introduce related work in two topics: program reduction and LLMs for software engineering.
6.1. Program Reduction
DDMin (Zeller and Hildebrandt, 2002) initiated the research topic of program reduction. It treats the input as a list of elements, and consistently splits the list into halves. Then it iteratively attempts to reduce the input list by exploring subsets and their complements at varying levels of granularity, transitioning from coarse to fine. Hierarchical Delta Debugging (Misherghi and Su, 2006), short for HDD, parses the program input into a parsing tree, and performs DDMin on each level of the tree structure. Perses (Sun et al., 2018) avoids the generation of syntactically invalid program variants during reduction by formal syntax transformation. Vulcan, further pushes the limit of Perses via identifier/sub-tree replacement and local exhaustive search. All above works are not customized for certain languages, though having high generality, they lack semantic knowledge of a certain language for further reduction.
Besides, some tools are specifically designed for certain languages. C-Reduce (Regehr et al., 2012), incorporating various transformation passes for features in C, is the most effective reducer on this language. J-Reduce (Kalhauge and Palsberg, 2019, 2021) is a tool for Java bytecode reduction, it reformulates the bytecode reduction into dependency graph simplification. ddSMT (Niemetz and Biere, 2013) is designed for reducing programs in SMT-LIBv2 format. All these works leverage language features to reduce more effectively than language-generic tools.
Distinct from prior work, LPR synergistically combines LLMs and language-generic reduction tool to harness the advantages of both. Language-generic reducers stand out as their remarkable generality across multiple languages, while LLMs excel in further refining the programs with the domain knowledge on certain languages learned from large training sets. Language-specific tools typically demand considerable human effort to design and implement feature-related transformations for reduction, while LPR requires only a few lines of natural languages prompts, significantly reducing the effort involved.
6.2. LLMs for Software Engineering
Large Language Models (LLMs) have proved their remarkable capability of undertaking multiple text-processing tasks, including source code-related works. Recent works focus on applying LLMs to facilitate software engineering tasks, or assessing the effectiveness, potential and limitations of LLMs on software development and maintenance. Some research (Huang et al., 2023; Xia and Zhang, 2023; Xia et al., 2023b, a) focus on empirically applying LLMs on automatic program repair (APR). Huang et al. (Huang et al., 2023) performed an empirical study on improvement brought by model fine-tuning in APR. Xia et al. (Xia et al., 2023b) thoroughly evaluated 9 state-of-the-art LLMs across multiple datasets and programming languages, and demonstrated that directly applying LLMs has already significantly outperformed all existing APR techniques. Additionally, some works focus on LLMs’ performance w.r.t. code completion, generation and fuzzing (Liu et al., 2023; Zhong and Wang, 2023; Tian et al., 2023a; Deng et al., 2023a, b), by leveraging the code analysis and generation ability of LLMs.
Similar to these studies, our approach LPR leverages LLMs for a software engineering task, i.e., program reduction. LPR harnesses the comprehension and generation capabilities of LLMs to refine the results of program reduction. However, our work distinguishes itself in the nature of the programs processed by LLMs. In related research, programs are typically logical and goal-oriented, often designed to fulfill a specific purpose. In contrast, the programs involved in our program reduction task are random, chaotic, and lack a clear objective. Consequently, our research sheds light on the performance of LLMs when dealing with programs that do not have an easily discernible purpose.
7. Conclusion
This paper proposes LLMs-aided program reduction (LPR), which is the first approach that leverages LLMs for the program reduction task to the best of our knowledge. By combining the strength of LLMs and existing language-generic program reduction techniques, LPR can perform language-specific transformations to effectively reduce the program while being language-generic (i.e., can be easily applied to a wide range of languages). The evaluation shows that in 50 benchmarks across three programming languages. LPR significantly outperforms Vulcan. Specifically, LPR produces 24.93%, 4.47%, and 11.71% smaller programs on C, Rust, and JavaScript, respectively. Meanwhile, The evaluation also demonstrates that LPR complements Vulcan to some extent. By reducing the outputs of LPR with Vulcan, we attained results that have similar sizes to those of C-Reduce in Benchmark-C. In terms of efficiency, LPR excels in reducing complex programs, and takes %, %, % less time than Vulcan to finish all the benchmarks, respectively.